A shipment of five computers contains two that are slightly defective. if a retailer receives three of these computers at random, list the elements of the sample paces using the letters D and N for defective and non-defective computers,respectively. To each sample point assign a value x of the random variable X representing the number of computers purchased by the ratailer which are slightly defective.

1
Expert's answer
2022-02-15T17:19:50-0500

Sample space:


"S=\{NNN, NND,NDN,DNN, NDD,"




"DDN, DND\}"

Let "X=" the number of slightly defective computers. The possible values of "X" are

"0,1,2."



"P(X=0)=\dfrac{\dbinom{2}{0}\dbinom{5-2}{3-0}}{\dbinom{5}{3}}=\dfrac{1(1)}{10}=\dfrac{1}{10}"




"P(X=1)=\dfrac{\dbinom{2}{1}\dbinom{5-2}{3-1}}{\dbinom{5}{3}}=\dfrac{2(3)}{10}=\dfrac{3}{5}"

"P(X=2)=\dfrac{\dbinom{2}{2}\dbinom{5-2}{3-2}}{\dbinom{5}{3}}=\dfrac{1(3)}{10}=\dfrac{3}{10}"